OpenStudy (anonymous):

Working with arithmetic sequences: a1 = 11 last term = 151 = a? What is the formula to figure out the last term?

OpenStudy (shubhamsrg):

you need to know the common difference..so ques incomplete as of now..

OpenStudy (shubhamsrg):

altough formulla for nth term : a_n = a + (n-1)d where a is first term n is term no. n a_n is nth term and d is common difference..

OpenStudy (anonymous):

Well the overall question I'm trying to answer is: What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 24 ? I know how to find "d" using the formula for an arithmetic sequence, but since I do no know what exactly the nth term for 151 is in this case, I haven't been able to use the formula to get "d".

OpenStudy (anonymous):

hey u need d u find it

OpenStudy (anonymous):

So then how do I find d with the info that I have?

OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 24 ? \[a_6=a_1 + (6-1)d\] \[24 = 4 + 5d\] \[20=5d\] \[d=4\]

OpenStudy (anonymous):

Woops. posted the wrong question: What is the sum of a 36-term arithmetic sequence where the first term is 11 and the last term is 151?

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

RIGHT THEREEE, needed that 36

OpenStudy (anonymous):

a_n = 151 a_1 = 11 n = 36 can you find d?

OpenStudy (anonymous):

d = 4 And staying up too long = stupidity

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

yup, so you need to find the sum now? "What is the sum of a 36-term arithmetic sequence where the first term is 11 and the last term is 151?"

OpenStudy (anonymous):

S36 = 2916

OpenStudy (anonymous):

I think I've learned more on here than on any of my eschool lessons.

OpenStudy (anonymous):

haha thanks, kinda feel flattered. and you got it right

OpenStudy (anonymous):

Thanks for the help